# Flow to the Mode: Mode-Seeking Diffusion Autoencoders for State-of-the-Art Image Tokenization

### v1

@@ Appendix C Rate-distortion-perception tradeoff
It could be argued that the desired “mode-seeking” behavior could be accomplished by simply increasing the sample likelihood. Unfortunately, evaluating the sample likelihood requires solving the log-determinant of the jacobian of the flow field [35], making it impractical to optimize directly. Neglecting the impact of this term we may increase or decrease the likelihood of $x_{1}$ by modulating its norm, since it is a centered isotropic gaussian (somewhat analogous to truncation sampling in GANs [5]).
@@
We attempt this analysis below. Whereas our proposed shifted sampler and post-training strategy improve both PSNR and rFID simultaneously, the naive sampling strategy described above only allows us to trade off rFID and PSNR against each other, constrained by the “rate-distortion-perception tradeoff” [4]. Importantly, rFID and PSNR cannot both be arbitrarily increased. See Figure 9 for a quantitative study of this process. Where the $x_{1}$ likelihood is increased too high, the result is an eventual degradation in the sample’s perceptual quality, while PSNR continues to increase. Even if it were possible to maximize the sample likelihood, as noted in [53], directly maximizing sample likelihood is inappropriate for image realism.
@@
Figure 9: Analysis of FlowMo with respect to the rate-distortion-perception tradeoff [4]. Varying $x_{1}$ likelihood sweeps out different rFID/PSNR curves at different rates. For this study, we use 5,000 samples for all metrics to make it computationally feasible.

### v3

@@ C.1 Rate-distortion-perception tradeoff
It could be argued that the desired “mode-seeking” behavior could be accomplished by simply increasing the sample likelihood. Unfortunately, evaluating the sample likelihood requires solving the log-determinant of the jacobian of the flow field [flow_matching], making it impractical to optimize directly. Neglecting the impact of this term we may increase or decrease the likelihood of $x_{1}$ by modulating its norm, since it is a centered isotropic gaussian (somewhat analogous to truncation sampling in GANs [biggan]).
@@
We attempt this analysis below. Whereas our proposed shifted sampler and post-training strategy improve both PSNR and rFID simultaneously, the naive sampling strategy described above only allows us to trade off rFID and PSNR against each other, constrained by the “rate-distortion-perception tradeoff” [ratedistortionperception]. Importantly, rFID and PSNR cannot both be arbitrarily increased. See Figure 9 for a quantitative study of this process. Where the $x_{1}$ likelihood is increased too high, the result is an eventual degradation in the sample’s perceptual quality, while PSNR continues to increase. Even if it were possible to maximize the sample likelihood, as noted in [universalcritic], directly maximizing sample likelihood is inappropriate for image realism.
@@
Figure 9: Analysis of FlowMo with respect to the rate-distortion-perception tradeoff [ratedistortionperception]. Varying $x_{1}$ likelihood sweeps out different rFID/PSNR curves at different rates. For this study, we use 5,000 samples for all metrics to make it computationally feasible.
C.2 Dataset and resolution generalization.
FlowMo-Lo is a diffusion autoencoder, so it naturally generalizes to higher resolutions by employing a patchwise diffuse-and-blend strategy known in prior work [highfidelity]. FlowMo-Lo achieves strong performance on unseen data, at 256 and 512 resolution, in terms of rFID, despite a lower BPP and only training on ImageNet-1K at $256\times 256$. In Table 8, we compare with the Cosmos DI-8x8 tokenizer, a generalizable tokenizer trained at multiple resolutions on a large internet-scale dataset.
ImageNet-1K OpenImages Food-101
Resolution 256 512 256 512 256 512
Cosmos DI-8x8 0.87 0.39 0.85 0.52 1.05 0.94
FlowMo-Hi 0.55 0.30 0.69 0.40 1.39 0.68
Table 8: Multiple resolutions. FlowMo can perform encoding and decoding at arbitrary resolutions following a patchwise diffusion strategy known in prior work [highfidelity].
C.3 Tokenizer scaling.
FlowMo is a transformer-only architecture parameterized in $\mu P$ to facilitate easy scaling. We show in the table below that, holding all else constant, all metrics improve every time the width factor is increased. For this experiment we train for 200K steps only. Other config settings are the same as FlowMo (fewer params).
$\mu P$ width # Params ($\times 10^{6})$ rFID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
2 260 7.77 20.84 0.169
3 367 5.31 21.28 0.160
4 517 4.45 21.60 0.155
5 710 3.84 21.70 0.152
C.4 “Mode-seeking,” precision and recall
We illustrate the ‘mode-seeking’ effect of post-training below. We compare the statistics of decompressed images to reference dataset statistics via the Precision and Recall metrics [precrec]. When there is a 1:1 correspondence between the elements of the conditioning dataset and the reference dataset, post-training improves both precision and recall. This corresponds to reducing the diversity of the distribution $p(x|c)$ for each $c$ by concentrating around its modes, which improves fidelity to the original reference dataset.
By contrast, when comparing to the train set statistics (i.e. an unrelated set of images not in 1:1 correspondence with the conditioning dataset), post-training improves precision (quality) at the cost of recall (diversity).
vs. train stats. vs. val stats.
Model name $uparrow Prec.\uparrow Rec.\uparrow$
FlowMo (fewer params) (no posttrain) 0.734 0.660 0.974 0.988
FlowMo (fewer params) 0.766 0.634 0.993 0.991
C.5 More results with post-training
The post-training scheme proposed generalizes to multiple architectures, consistently improving their performance. We have added more results extending Table 5 below, for FlowMo (fewer params), which is defined in Table 6, and FlowMo-Continuous, which is the version against which we compare DiTo in Table 3.
Model name rFID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
FlowMo (fewer params) (no post-train) 2.02 21.32 0.143
FlowMo (fewer params) 1.21 22.20 0.120
FlowMo-Continuous (no post-train) 0.74 25.65 0.064
FlowMo-Continuous 0.65 26.61 0.054
C.6 Wall clock times
We provide wall-clock times below. Note that FlowMo encoding is comparable to LlamaGen-32, so it does not bottleneck second-stage generative modeling. The main slowdown is due to a diffusion decoder, which we consider acceptable because it enables significantly better reconstruction. Decoding could be made faster ($\approx$ $5\times$ to $10\times$) via distillation.
Model name Encode 1 image (s) Decode 1 image (s)
LlamaGen-32 0.021 0.038
FlowMo-B 0.050 2.391
FlowMo-B (small) 0.011 0.322